OpenStudy (usgciceman):

Which line set of points are part of a line that is parallel to the line y = (-⅔)x ­ + 1 (5, 0), (­-1, 4) (2, 0), (2, -­1) (2, ­-2), (6, 4) (3, -­2), (6, 0)

OpenStudy (anonymous):

A parallel line has to have the same gradient, so work out the gradient for each set of values, the gradient of a parallel line will be... -2/3 Eg. For the first set of values... \[Gradient =\frac{ 5- (-1) }{ 0-4}=\frac{ 6 }{ -4 }=\frac{ 3 }{ -2 }\]